Two numbers sum to give an answer of 120. One number is 48 greater than the other number. Find both numbers.

Answer by Maths68(1474)

Let x be the greater and y be the lesser number

Two numbers sum to give an answer of 120
x+y=120................(1)

One number is 48 greater than the other number
x=y+48..................(2)
Put the value of x from (2) to (1)
y+48+y=120
2y+48=120
2y=120-48
2y=120-48
2y=72
Divide by 2 both sides
2y/2=72/2
y=36
Put the value of y in (2)
x=y+48..................(2)
x=36+48
x=84
So numbers are 84 and 36